# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v3

@@ 2. $(2)$
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
Once we know $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, it remains to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, where $I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b+1,j_{1}b\right]$ and $I_{j_{1}}^{\prime}=\left[p_{j_{2}-1}b+1,j_{2}b\right]$. Since $j_{2}-j_{1}\geq 2$, we have $(p_{j_{2}-1}-1)b+2>p_{j_{1}}b$. This implies that the $b$-burst-deletion in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ does not affect $\mathbf{x}_{I_{j_{1}}^{\prime}}$. Therefore, we conclude that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So to decode $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}$, we need the function $\phi(\cdot)$ defined in Lemma II.2.
Next, we will apply $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ to decode $\mathbf{x}_{I_{j}^{\prime}}$. Recall that $U(\mathbf{x})_{1}$ is $d$-regular. So by definition, we have $\left|I_{j}^{\prime}\right|=(p_{j}-p_{j-1})b\leq db\log n$. Then by Lemma II.2, $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{\log(db\log n)+O_{q,b}(1)}$. Let $Q_{1}\geq\max\{n/b,N_{2}\}$ be the smallest prime. Define
$g^{(0)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{2N_{2}},$ (15)
$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{1}}.$
@@
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Furthermore, suppose that $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known. Then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}$.
@@
Since $j_{2}-j_{1}\geq 2$, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So by Lemma II.2, to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, it it sufficient to obtain $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known, it is easy to see that $\mathbf{x}_{I_{j}^{\prime}}$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{1}}$. Then it follows from Equation 15 that
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{2N_{2}},$ (16)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$ (17)
Since $0\leq\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right),\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 16 that $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)=\delta^{(0)}$. Combining this with Equation 17, we get
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{1}},$ (18)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$
Now similar to Lemma IV.6, we can uniquely solve $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$ from (18). ∎
V-B When Case (2) in Lemma V.2 happens
In this case, we can find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$. This implies that $\mathbf{x}_{k}$ is known to us for all $k\notin J$. Suppose $J=[c_{1},c_{2}]$. Then $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}\in\mathcal{B}_{b}^{2}\left(\mathbf{x}_{J}\right)$. To recover $\mathbf{x}_{J}$, we apply a trivial $2$-$b$-burst-deletion correcting code of length $2\left|J\right|$, which is given in Lemma II.5.
Let $P=3bd\log(n/b)+b-1$ and
@@
For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$. Let the function $\psi(\cdot)$ be defined in Lemma II.5. Since $\left|J_{j}\right|\leq 2P$, we can view each $\psi\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7b\left\lceil\log q\right\rceil\log(2P/b)+O_{q,b}(1)}$ and define
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{0}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{1}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}}.\end{array}$ (19)
@@
Suppose that we found an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$. Then given $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$, we can uniquely recover $\mathbf{x}_{J}$ (and thus $\mathbf{x}$) from $\mathbf{x}^{\prime}$.
Proof:
We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$. Without loss of generality, we assume $j_{0}\in L_{0}$. Then $\psi\left(\mathbf{x}_{J_{j}}\right)$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$. In addition, if $J_{j_{0}}=[c_{1},c_{2}]$, it is clear that $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}=\mathcal{B}_{2}^{b}\left(\mathbf{x}_{J_{0}}\right)$. It follows from Equation 19, that
$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, we have
$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
Now by Lemma II.5, one can uniquely recover $\mathbf{x}_{J_{0}}$ and thus $\mathbf{x}_{J}$. ∎
@@
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{2}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
@@
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{2}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma V.2, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemmas V.3 and V.4 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma V.5 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{2}$ can correct two $b$-burst-deletions.
By Lemma V.1, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $U(\mathbf{x})_{1}$ is $d$-regular. Now the claim about redundancy follows from the pigeonhole principle. ∎
Remark V.2
Let $q=2$. Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $6\log n+(14b+11)\log\log n+O_{b}(1)$. Comparing this with Theorem IV.1, we can see that when $2\leq b\leq 16$, the code in Theorem IV.1 has lower redundancy, while when $b\geq 17$, the code in Theorem V.1 has lower redundancy.
VI Non-binary Two-deletion Correcting Codes
In this section, we always assume $q>2$. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.4 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23]
Lemma VI.1
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma VI.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. Define
The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ will help to deal with case (1) in Lemma IV.4. For case (2) in Lemma IV.4, let $P=3d\log n$ and
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{0}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{1}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}.\end{array}$ (21)
Now we are ready to give our code.
Theorem VI.1
Suppose $q>2$. Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{1},c_{2}<N_{2}$, define the code $\mathcal{C}_{3}$ as
$\mathcal{C}_{3}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h_{e}(\mathbf{x})=c_{1},h_{o}(\mathbf{x})=c_{2}\end{array}\right\}$
Then $\mathcal{C}_{3}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{1}$ and $c_{2}$, such that the redundancy of $\mathcal{C}_{3}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
  • $q$ is even, $d=7$ and $n\geq 9$;
The proof of Theorem VI.1 is similar to those of Theorems IV.1 and V.1, and thus omitted.
VII Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $7\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. Furthermore, when $b=2$, we show that there is a binary code with redundancy at most $6\log n+O\left(\log\log n\right)$. Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.

### v4

@@ V Non-binary Two-deletion Correcting Codes
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma V.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. By Lemma II.6, we have $n\leq Q<2n$. Define
The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ can help to deal with case (1) in Lemma IV.3, as shown in the next lemma.
@@
Suppose that the two deletions in $\mathbf{x}$ occurred in two known intervals $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$, respectively. Then given $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
@@
Notice that $x_{k}$ is known to us for all $k\notin I_{j_{1}}\cup I_{j_{2}}$. It remains to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$. It is easy to see that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}+1,p_{j_{1}}-1\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2},p_{j_{2}}-2\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{2}}}\right)$. By Lemma V.1, to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$, it is sufficient to know the values of $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)$. Note that $\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)$ is known to us for all $j\neq j_{1},j_{2}$. Let $\delta_{0}=\left(f_{0}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{2N_{1}}$ and $\delta_{1}=\left(f_{1}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{Q}$. Then it follows from Equation 18 that
$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{2N_{1}},$ (19)
$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}.$ (20)
Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<2N_{1}$ and $0\leq\delta_{0}<2N_{1}$, we can conclude from Equation 19 that $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\delta_{0}$. Combining this with Equation 20, we get
$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{Q},$ (21)
$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}.$
Since $1\leq j_{1}<j_{2}\leq n\leq Q$, we have $j_{1}\not\equiv j_{2}\pmod{Q}$. Therefore, system (21) has a unique solution in the field $\mathbb{F}_{Q}$: $\left(\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q},\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}\right)$. Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right),\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<N_{1}\leq Q$, we have $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q}$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}$. Now the proof is completed. ∎
For case (2) in Lemma IV.3, we follow similar idea in Section IV-C. Let $P=3d\log n$ and
@@
$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{s}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}.\$ (22)
@@
Suppose we are in case (2) in Lemma IV.3. Given $h^{(0)}$ and $h^{(1)}$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
The proof is similar to that of Lemma IV.7, and thus omitted.
Our $q$-ary two-deletion correcting code is given in the next theorem.
@@
Suppose $q>2$. Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
Then $\mathcal{C}_{2}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
@@
The proof of Theorem V.1 follows directly from Lemmas V.2 and V.3. The claim for redundancy follows from Lemma IV.4 and Lemma II.6.
VI Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $5\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.